# Are We in the AI-Generated Text World Already? Quantifying and Monitoring AIGT on Social Media

### v1

@@ Appendix D Introduction of Open Source Datasets for Training Detectors
  • Given that high-quality Supervised Finetuning (SFT) datasets are frequently used for finetuning LLMs, and considering the lack of Claude and GPT-4 model-related content in the AIGT detection datasets, we also incorporate four SFT datasets with instruction-following features: Claude2-Alpaca [^2] https://github.com/Lichang-Chen/claude2-alpaca, Claude-3-Opus-Claude-3.5-Sonnet-9k [^3] https://huggingface.co/datasets/QuietImpostor/Claude-3-Opus-Claude-3.5-Sonnnet-9k, GPTeacher/GPT-4 General-Instruct [^4] https://github.com/teknium1/GPTeacher/tree/main/Instruct, and Instruction in the Wild [^5] https://github.com/XueFuzhao/InstructionWild.
@@ 6. Respond with the revised content only and nothing else:
Appendix G Details About the Collection of High-Frequency Words and Model Interpretation Analysis Methods
@@
We use the Spacy library [15] to classify the part-of-speech of words in the SM-D, specifically dividing them into adjectives, adverbs, and connectives. We then select around the top 20 words for human-preferred and AI-preferred categories, respectively. For detailed results, refer to Table A4.
@@
Dataset Type Sentence Number
MGT-Academic [23] Llama3 $1,478,485$
Mixtral 8$\times$7B $2,639,498$
Moonshot $726,357$
GPT-3.5 $1,611,244$
Human $6,007,476$
Coco-GPT3.5 [21] GPT-3.5 $79,647$
Human $55,565$
GPABench2 [24] GPT-3.5 $12,648,338$ (Sample)
Human $1,065,860$
LWD [40] Llama2 $94,732$
GPT-3.5 $95,443$
GPT-4 $62,632$
Human $106,952$
AIGT [38] Llama2 $6,967$
Alpaca 7B $6,083$
Vicuna 13B $7,028$
GPT-3.5 $8,022$
GPT-4 $7,156$
Human $12,228$
HC3 [12] GPT-3.5 $184,692$
Human $347,423$
Table A1: Statistics of open-source datasets (Part 1).
Dataset Type Sentence Number
Claude2-Alpaca Claude-2 $404,051$
Claude-3-Opus-Claude-3.5-Sonnnet-9k Claude-3 $276,246$
Human $37,785$
GPTeacher/GPT-4 General-Instruct GPT-4 $74,160$
Human $24,465$
Alpaca_GPT4 GPT-4 $354,801$
Human $22,253$
Instruction in the Wild GPT-3.5 $300,424$
Table A2: Statistics of open-source datasets (Part 2).
Figure A1: Token length distribution in the training set, calculated by the Llama-2 tokenizer [45].
Figure A3: Timeline of authors’ earliest adoption of AIGTs.
@@
Table A3: Sentence number statistics of our generated datasets (Llama Series include Llama-1, 2, 3; GPT Series include GPT-3.5, GPT-4o-mini).

### v2

@@ Conclusion
  4. 4.
    Inability of the AIGT detector to detect texts generated by early models release before ChatGPT: During the dataset collection and training process, we only included models that were released after November 2022. This decision was made because our study specifically focuses on more powerful models such as ChatGPT. We acknowledge that this approach may lead to misclassifications for earlier models; however, this does not affect the overall experimental results or the validity of our findings.
@@ Appendix D Introduction of Open Source Datasets for Training Detectors
  • Given that high-quality Supervised Finetuning (SFT) datasets are frequently used for finetuning LLMs, and considering the lack of Claude and GPT-4 model-related content in the AIGT detection datasets, we also incorporate four SFT datasets with instruction-following features: Claude2-Alpaca [^1] https://github.com/Lichang-Chen/claude2-alpaca., Claude-3-Opus-Claude-3.5-Sonnet-9k [^2] https://huggingface.co/datasets/QuietImpostor/Claude-3-Opus-Claude-3.5-Sonnnet-9k., GPTeacher/GPT-4 General-Instruct [^3] https://github.com/teknium1/GPTeacher/tree/main/Instruct., and Instruction in the Wild [^4] https://github.com/XueFuzhao/InstructionWild..
@@ 6. Respond with the revised content only and nothing else:
Appendix G Detailed Performance of OSM-Det
Table A6 presents the performance of OSM-Det on individual platform-specific datasets within AIGTBench. The results show that OSM-Det achieves consistently high accuracy across all three platforms, with accuracy scores of $0.995$, $0.999$, and $0.984$ on Medium, Quora, and Reddit, respectively.
Figure A7 illustrates the accuracy and F1-score across different text lengths in AIGTBench. We observe that accuracy is relatively lower for shorter texts, with an accuracy of approximately $0.940$ for texts between $0-149$ characters. However, for texts exceeding $150$ characters, accuracy improves significantly to above $0.980$. Both accuracy and F1-score continue to increase as text length increases.
To ensure the reliability of our conclusions, we filter out texts shorter than $150$ characters from SM-D in our paper.
Appendix H Generalizability of OSM-Det
AIGTBench is a comprehensive dataset that contains multi-source, multi-domain, and multi-LLM data. The diversity of this dataset enhances the generalizability of detectors in real-world (in-the-wild) environments. OSM-Det, on the other hand, is the optimal detection model trained on AIGTBench.
In this section, we evaluate the generalizability of OSM-Det from three perspectives: AIGTs produced with different generation parameters, AIGTs of social media generated by unseen models, and tests in the wild.
Different Generation Parameters. To investigate whether OSM-Det can effectively detect AIGTs generated with different generation parameters, we randomly sample $5,000$ HWTs from the AIGTBench and apply the same prompt to refine them using different generation parameters (including temperature, top-p, and top-k). The models used for this experiment are GPT4o and GPT4o-mini.
As shown in Figure 2, OSM-Det maintains an accuracy of over $0.99$ across the entire range of temperature settings (0.1 to 1.0). Top-P (0.1 to 1.0) and Top-K (1 to 200) show a similar trend. This indicates that OSM-Det demonstrates strong generalizability when detecting AIGTs generated with different parameters.
AIGTs of Social Media Generated By Unseen Models. To investigate the generalizability of OSM-Det on social media AIGTs generated by unseen models, we selected $6$ pre-trained models, including Deepseek-V3 [25], GLM-4-Flash [5], Gemini-1.5-Flash [9], Gemini-2.0-Flash [9], Yi-1.5-34B [64], and InternVL2.5-8B [42]. Additionally, we include three fine-tuned models based on the LLaMA series: Dolphin3.0-Llama3.1-8B [8], Llama3-OpenBioLLM-8B [1], and Xwin-LM-13B-V0.2 [63]. Since none of these models were included in AIGTBench, they are considered unseen to OSM-Det. We also apply the same polishing process to the previously selected $5,000$ data samples.
From Table A8, we observe that OSM-Det maintains strong detection performance across these unseen models. The lowest performance was recorded for InternVL2.5-8B, yet it still achieve an accuracy of $0.925$ and an F1-score of $0.958$. This demonstrates that OSM-Det exhibits strong generalization capability when detecting AIGTs generated by previously unseen LLMs.
Test In the Wild. To test OSM-Det in the wild, we randomly select datasets from the huggingface platform for evaluation. These datasets are in two main categories: unseen models and unseen domains, neither of which are included in AIGTBench.
As shown in Table A7, for the unseen model scenario, the test results align with previous findings, where OSM-Det maintains high accuracy and F1-score. Similarly, in the unseen domain scenario, OSM-Det also demonstrates strong generalizability, achieving a minimum accuracy of $0.943$. This is consistent with the findings of Liu et al. [28], which suggest that AIGT detectors exhibit generalizability across different domains.
Appendix I Details About the Collection of High-Frequency Words and Model Interpretation Analysis Methods
@@
We use the Spacy library [19] to classify the part-of-speech of words in the AIGTBench, specifically dividing them into adjectives, adverbs, and connectives. We then select around the top 20 words for human-preferred and AI-preferred categories, respectively. For detailed results, refer to Table A5.
@@
Table A1: Sentence number statistics of our generated datasets (Llama Series include Llama-1, 2, 3; GPT Series include GPT-3.5, GPT4o-mini).
Figure A1: Proportion of total sentences various LLMs, with “Others” including Alpaca 7B and Vicuna 13B.
(a) Token length distribution in the training set.
(b) Token length distribution in the testing set.
(c) Token length distribution in the validation set.
Figure A2: Token length distribution in the training, testing, and validation sets, calculated by the Llama-2 tokenizer [57].
Platform # text (Human) FPR
Medium $116,303$ $1.82\%$
Quora $101,145$ $1.36\%$
Reddit $53,321$ $1.70\%$
Table A2: FPR of OSM-Det on social media platforms.
(a) Number of Likes.
(b) Number of Comments.
Figure A4: Differences between predicted AIGTs and predicted HWTs compressed using a log10 transformation.
Figure A5: Word frequency trends on Reddit from January 1, 2022, to July 31, 2024.
Figure A6: Timeline of authors’ earliest adoption of AIGTs.
Dataset Type Sentence Number Domain
MGT-Academic [28] Llama3 $1,478,485$ STEM (Physics, Math, Biology, CS, EE, Statistics, Chemistry, Medicine), Social Science (Education, Economy, Management), Humanities (Literature, Law, Art, History, Philosophy)
Mixtral 8$\times$7B $2,639,498$
Moonshot $726,357$
GPT-3.5 $1,611,244$
Human $6,007,476$
Coco-GPT3.5 [26] GPT-3.5 $79,647$ News
Human $55,565$
GPABench2 [29] GPT-3.5 $12,648,338$ (Sample) Computer Science, Physics, Social Sciences
Human $1,065,860$
LWD [52] Llama2 $94,732$ Finance, Social Media
GPT-3.5 $95,443$
GPT-4 $62,632$
Human $106,952$
AIGT [50] Llama2 $6,967$ Soical media, News, Academic Writing
Alpaca 7B $6,083$
Vicuna 13B $7,028$
GPT-3.5 $8,022$
GPT-4 $7,156$
Human $12,228$
HC3 [15] GPT-3.5 $184,692$ Open-domain, Finance, Medicine, Law, and Psychology
Human $347,423$
Table A3: Statistics of open-source datasets (part 1).
Dataset Type Sentence Number Domain
Claude2-Alpaca Claude-2 $404,051$ Open-domain
Claude-3-Opus-Claude-3.5-Sonnnet-9k Claude-3 $276,246$ Open-domain
Human $37,785$
GPTeacher/GPT-4 General-Instruct GPT-4 $74,160$ Open-domain
Human $24,465$
Alpaca_GPT4 GPT-4 $354,801$ Open-domain
Human $22,253$
Instruction in the Wild GPT-3.5 $300,424$ Open-domain
Table A4: Statistics of open-source datasets (part 2).
@@
Platform Accuracy F1-score
Medium $0.995$ $0.995$
Quora $0.999$ $0.999$
Reddit $0.984$ $0.984$
Table A6: Performance of OSM-Det on AIGTs within AIGTBench across different platforms.
Figure A7: Performance of OSM-Det across varying text lengths on AIGTBench.
Category Dataset Performance
Accuracy F1-score
Unseen Model QwQ-32B-Preview [31] 0.999 0.999
Gemini-2.0-Flash [44] 0.993 0.997
Deepseek-R1-Llama-70B [32] 0.999 0.999
Unseen Domain Roleplay-English [38] 0.999 0.999
Mannerstral-dataset [18] 0.943 0.968
InternVL-SA-1B-Captio [43] 0.998 0.999
Table A7: Test OSM-Det in the wild. (all datasets from HuggingFace)
Model Accuracy F1-score
Deepseek-V3 0.986 0.993
GLM-4-Flash 0.997 0.998
Gemini-1.5-Flash 0.938 0.952
Gemini-2.0-Flash 0.984 0.992
Yi-1.5-34B 0.999 0.999
InternVL2.5-8B 0.925 0.958
Dolphin3.0-Llama3.1-8B 0.996 0.998
Llama3-OpenBioLLM-8B 0.960 0.980
Xwin-LM-13B-V0.2 0.996 0.998
Table A8: Performance of OSM-Det on AIGTs generated from unseen LLMs based on social media data from AIGTBench.